# On the Client Preference of LLM Fine-tuning in Federated Learning

### v1

@@ 3.2.2 Algorithm Design
3.3 Discussion
We discuss the limitations of FedBis which motivate us to propose FedBiscuit.
Preference Heterogeneity.
A performance gap between FedBis and centralized training could arise from data heterogeneity among clients, a common issue in FL. Different from centralized training that aggregates all the clients’ data and samples an i.i.d. batch in each training round, FedBis samples a subset of clients in each round, with each client independently optimizing the model based on their local data. This could result in a global aggregation that diverges from the global optimum Karimireddy et al. (2020); Wu et al. (2023).
Reward Hacking.
As demonstrated in experiments, FedBis’s performance improves first but may later decline with the increase of training rounds. This phenomenon, known as reward hacking, is discussed by Skalse et al. (2022) as an inevitable issue in training a reward proxy model, which is used to enhance the performance of a policy model (e.g., LLM). However, we can mitigate this impact by delaying the inflection point, allowing the reward proxy model to continue improving performance for more training rounds and ultimately achieve a higher rating.
4 FedBiscuit: FedBis with Cluster-wise Aggregation
In this section, we aim to address the aforementioned limitations of FedBis. To tackle reward hacking, Eisenstein et al. (2023) and Coste et al. (2024) introduce a promising approach that trains multiple reward models at the same time because aggregation over multiple reward model outputs can provide a more robust reward estimate. Furthermore, recognizing that some clients may share similar preferences, we employ clustered FL Sattler et al. (2020); Ghosh et al. (2020); Ma et al. (2023) to group clients with similar preferences for joint model training. Notably, these two approaches complement each other, inspiring us to combine them into a novel algorithm FedBiscuit that simultaneously combats reward hacking and preference heterogeneity.
Problem Formulation.
In this work, we consider training multiple binary selectors of $U$. To ensure that all selectors are trained without bias towards a small specific group, we mandate that these selectors be trained using evenly disjoint clusters of clients. Additionally, a client’s preference should align more closely with those within the same cluster than with those in different clusters. To this end, we can formulate the following objective:
$\min_{\phi_{[U]}\in\mathbb{R}^{U\times d}}F({\phi_{[U]}})\overset{\triangle}{=}\sum_{m\in[M]}p_{m}\left(\min_{u\in[U]}F_{m}({\phi_{u}})\right)$
$s.t.\quad\max\{|M_{u}|\}_{u\in[U]}-\min\{|M_{u}|\}_{u\in[U]}\leq 1$ (7)
where the function $F_{m}$ follows the same definition of Equation (2). $\phi_{u}$ indicates the $u$-th binary selector, and $M_{u}$ means a set of clients using the $u$-th selector. By definition, $\cup_{u\in[U]}M_{u}=[M]$, and $\cap_{u\in[U]}M_{u}=\emptyset$.
Next we explore how the proposed FedBiscuit optimizes Equation (7).
4.1 Algorithm Design
Section 3.1 mentions that a client $m\in[M]$ holds a preference dataset $\mathcal{D}_{m}$. Before the model training, client $m$ splits her dataset into two disjoint sets, namely, a training set $\mathcal{D}_{m,train}$ and a validation set $\mathcal{D}_{m,val}$, where $|\mathcal{D}_{m,train}|>>|\mathcal{D}_{m,val}|$.
The proposed FedBiscuit consists of two phases: 1) We train each selector for a couple of rounds so that all $U$ selectors have fundamental capacities in selecting the preferred completion, and 2) we divide the clients into disjoint clusters of size $U$ and train each binary selector with a specific cluster.
Phase 1: Warm-up.
In the beginning, we initialize each binary selector $\phi_{u}(u\in[U])$ with an identical pretrained LLM. Subsequently, starting from $u=0$, we train a selector $\phi_{u}$ for $R_{pre}$ consecutive communication rounds following the steps of FedBis: In each communication round, the server samples a subset of client $\mathcal{A}$ and broadcasts the selector $\phi_{u}$ to them. Each client $m\in\mathcal{A}$ then locally trains the selector for $K$ iterations using the dataset $\mathcal{D}_{m,train}$. At the end of the communication round, the server aggregates and updates the selector $\phi_{u}$ via Equation (4). After completing the training of $\phi_{u}$, the server initiates the training of the next selector $\phi_{u+1}$ by repeating the above steps until all selectors are trained.
Phase 2: Clustered FL Training.
After the first phase, we obtain $U$ different selectors, denoted as $\{\phi_{u,0}\}_{u\in[U]}$. Unlike FedBis, this phase includes an additional step called client grouping, which partitions the clients into multiple disjoint clusters based on their preferences. In each communication round $r\in[R]$, the proposed FedBiscuit optimizes all the selectors $\phi_{[U]}$ using the following four steps:
Step 2.1: Client Grouping. This step is executed every $\tau$ communication rounds, i.e., when $r$ can be divided by $\tau$, or $\tau|r$. During this step, the server broadcasts all selectors $\phi_{[U],r}$ to all clients $[M]$. Then, a client $m$ calculates the averaged loss for each selector $\phi_{u,r}$ using local validation set via $\frac{1}{|\mathcal{D}_{m,val}|}\sum_{(x,y_{0},y_{1},i)\sim\mathcal{D}_{m,val}}\left[\ell_{CE}(i|\phi_{u,r};x,y_{0},y_{1})\right]$. The server thereby collects all these losses and adopts a greedy clustering approach Sattler et al. (2020); Ma et al. (2023) to assign each client to the selector where they achieve the minimum loss. However, an obvious deficiency is an imbalance where some selectors are chosen by many clients and others by few. It is noted that the selectors trained with more clients achieve remarkable performance, while some may be overfitted to a specific group of clients. Therefore, the greedy clustering approach negatively impacts the overall performance when building a global preference dataset. To tackle the limitation, we propose to balance the clusters using the following steps repeatedly until the clients are evenly distributed:
  • Choose the cluster selected by the most clients.
  • If the cluster can accommodate $n$ clients, cap the cluster at $n$ clients and reassign the rest to other clusters where they achieve suboptimal loss.
Finally, we obtain balanced and disjoint clusters. Let a client $m$ train with the $U_{m}$-th selector $\phi_{U_{m}}$ for the next $\tau$ rounds. After client grouping step, the proposed method proceeds to the following three steps as outlined in FedBis.
Step 2.3: Local Training. The client $m\in\mathcal{A}$ receives a binary selector $\phi_{U_{m},r}$ from the server and trains the selector for $K$ iterations following the update rule of Equation (3). Finally, let the updated local selector be $\phi_{U_{m},r,K}^{m}$, and the client pushes it to the server.
Step 2.4: Global Aggregation. The server collects updated selectors from all participants $\mathcal{A}$. Since there are several binary selectors, the server updates each one with a designated group of clients intended to train on that specific selector. For instance, the aggregation rule for the selector $u\in[U]$ follows
$\phi_{u,r+1}=\left(1-\sum_{m\in\mathcal{A}_{u}}p_{m}\right)\phi_{u,r}+\sum_{m\in\mathcal{A}_{u}}p_{m}\phi_{u,r,K}^{m}$ (8)
It is noted that performance degradation occurs when a model is trained by clients with time-varying sizes in FedAvg Gu et al. (2021); Wang and Ji (2023). In other words, Equation (4) is no longer suitable for multi-selector aggregation due to the fluctuation in the number of clients training a selector in each communication round. Therefore, FedBiscuit adopts a new aggregation rule as formulated in Equation (8).
FedBiscuit finally produces a set of well-trained selectors $\phi_{[U],R}$ and he subsequent objective is to enhance LLM performance with the help of these selectors, as explored below.
Reinforcement-learning Fine-tuning with Multiple Selectors.
We can leverage the methodology mentioned in Section 3.2, and one of the key steps involves constructing a preference dataset incorporating multiple selectors. For this, we employ a strategy of majority voting. Given an instruction $x\in\hat{\mathcal{D}}$ and a pair of generated completions $(y_{0},y_{1})$, we assume a selector $u\in[U]$ prefers $y_{i_{u}}$, where $i_{u}\in\{0,1\}$. Therefore, the pair is assigned a label $i=\arg\max\{i_{u}\}_{u\in[U]}$, meaning that the completion $y_{i}$ is favored by the most clients.
4.2 Discussion: Integration with LoRA
5 Federated Human Preference Benchmark
This section mainly focuses on how we prepare federated human preference datasets, while the next section introduces the experimental setup and analyzes numerical results. Specifically, we cover two of the most common NLP tasks, i.e., summarization and question-answering. All two datasets are partitioned based on the public datasets, and the following subsections will include the details. We will release these datasets on HuggingFace soon.
Summarization.
Stiennon et al. (2020) introduces a summarization dataset that consists of Reddit posts with human-written tl;dr Völske et al. (2017). This dataset consists of two parts, one is a pretrained dataset, while the other is a dataset with human preference. As suggested by Ouyang et al. (2022), we ensure a post does not appear in both datasets. We assume the pretrained dataset is stored on the server side, and 60% of data are served for model pertaining such that the model can perform well on summarization. The remaining 40% are used for the RLHF process to improve the LLM performance and generate human-preferred content. Since the human-preference dataset contains the worker ID, we partition the dataset based on the worker ID so that the dataset can be partitioned into 53 workers.
Question-Answering (QA).
We reconstruct the public dataset SHP, which comprises numerous questions from Reddit posts and their corresponding user answers Ethayarajh et al. (2022). The preference indicator is based on the number of likes an answer receives. Given that the dataset spans 18 domains, we partition the dataset using a Dirichlet distribution with a parameter of 0.3, ensuring that no questions overlap between clients. In our experiment, we prepare 300 clients, and Figure 2 visualizes the data distribution on the selected clients. For the RLHF process, we use a set of 2.6K Reddit questions.
6 Experiments
6.1 Experimental Setup
Model and computation environment.
We initialize the binary selector(s) using the pretrained LLaMA-2-7B Touvron et al. (2023), configuring the final layer to produce binary outputs "A" and "B" only. The LLM chosen for content generation depends on the tasks: (i) For the summarization task, we start with LLaMA-2-7B and fine-tune it using a pretrained dataset; (ii) For the QA task, we initialize the LLM with Alpaca-7B Taori et al. (2023). To reduce computation efforts, we employ LoRA to fine-tune the models. Our implementation, built upon FederatedScope Xie et al. (2023); Kuang et al. (2023), will soon be available on GitHub. The experiments are conducted on machines equipped with two Nvidia A100 GPU cards, Intel Xeon Platinum 8369B CPUs, and 256GB RAM.
Evaluation.
We evaluate two models produced by our proposed FedBis and FedBiscuit: a binary selector and an LLM. We employ different strategies to assess each model:
  • Binary selector: The evaluation includes two metrics: agreement and best-of-$n$. Agreement measures the hit rate of a selector against a preference dataset annotated by humans or ChatGPT. Additionally, we use the best-of-$n$ approach by selecting the best completion from $n$ generated by a task-specific LLM. We then evaluate the average rating for the selector’s choices using Auto-J Li et al. (2023a).
  • LLM: After reinforcement-learning fine-tuning, the LLM is evaluated on its ability to generate human-preferred content. This means we can assess the quality of the generated texts. For example, given an instruction set, the LLM produces one completion per instruction, and Auto-J evaluates the average rating of these completions. Furthermore, we compare the generated completions with a reference set of responses, annotated by humans or ChatGPT, and calculate a win rate based on how often the generated response is superior to the reference one.
Due to the space limit, the hyperparameter settings are presented in Appendix A. Moreover, Appendix B provides real cases to demonstrate the performance of the proposed FedBis and FedBiscuit. In particular, Appendix B.1 discusses the results under the QA task.
6.2 Numerical Results on Summarization
Selector LLM
Agreement Best-of-$n$ Rating Win Rate
SFT - - 5.028 29.71%
Centralized 73.10% 5.302 5.688 78.89%
FedBis 70.44% 5.274 5.661 71.35%
FedBiscuit 70.52% 5.305 5.703 80.65%
Table 1: Performce under summarization task. All values here indicate their best performance within 500 communication rounds of training.
In this section, the evaluation data originates from the TL;DR dataset, as mentioned in Section 5. The dataset comprises two disjoint parts: one ranked by a group of labelers for model-generated responses, and the other written by users to summarize the key content of a post. We use the former to compute the consistency between the selector and the human annotator. For the latter, we apply various metrics, including best-of-$n$, rating, and win rate. The results are presented in Table 1 and Figure 3.
The table shows that conventional centralized training outperforms the proposed FedBiscuit in terms of agreement. This is because the agreement evaluation data have a similar distribution to the training dataset, as their outputs are generated from the same language models and labeled by the same group of labelers Stiennon et al. (2020). Consequently, centralized training performs better than the proposed FedBis and FedBiscuit, which are affected by data heterogeneity.
However, when evaluating the selectors with datasets generated by a supervised fine-tuning model, the proposed FedBiscuit slightly outperforms centralized training. These results suggest that a centrally trained selector performs poorly in terms of generalization and is prone to overfitting to a specific dataset distribution. In contrast, comparing FedBiscuit with FedBis, we find that FedBiscuit mitigates data heterogeneity and produces a more robust selection of completion pairs.
Figure 3: Auto-J rating of best-of-$n$ against communication rounds under summarization tasks.
Figure 3 illustrates the performance trend across communication rounds. As discussed in Section 3.3, training a binary selector can lead to reward hacking. For both centralized training and FedBis, which train a single selector, we observe an inflection point where the selector’s performance begins to decline. However, this inflection point has not yet appeared in FedBiscuit, allowing it to continuously improve and eventually surpass the best performance of centralized training. It is important to note that the warmup rounds are included in the communication rounds, which explains FedBiscuit’s initial poor performance.
7 Conclusion
In this work, we explore a feasible framework to achieve RLHF in FL. Specifically, we train a binary selector across different clients using their local preference datasets, and then use the well-trained selector to align an LLM with human preferences. We propose two approaches to enable selector training: FedBis and FedBiscuit. FedBis provides a framework to train a single selector, while FedBiscuit ensembles multiple selectors to more robustly simulate human preferences. With the proposed federated human preference datasets, we conduct empirical studies to validate our statements and demonstrate the superiority of FedBiscuit.
Ethics Statement
Limitations
One notable limitation of our work lies in the construction of the preference dataset, which relies solely on publicly available data rather than gathering information directly from real clients. By doing so, we miss out on the nuances and intricacies of individual preferences that can only be captured through firsthand data collection. As a result, our dataset may lack the depth and breadth necessary to fully comprehend the true heterogeneity of preferences among clients. Without access to authentic client data, we may inadvertently overlook important variations in preferences, potentially limiting the applicability and robustness of our findings.
Another limitation pertains to the use of a task-specific dataset rather than a more generalized one encompassing a broader spectrum of tasks. While task-specific datasets offer advantages such as focused analysis and tailored insights, they may also restrict the scope of our research and hinder its generalizability. By incorporating a more diverse range of tasks into our dataset, we could gain a more comprehensive understanding of clients’ preferences across various domains, thereby enhancing the versatility and validity of our findings.
Additionally, our work employs a binary selector that implicitly assumes one response is superior to another, overlooking scenarios where responses may exhibit similar levels of quality. This oversimplified approach fails to leverage valuable data that could provide valuable insights into subtle differences and nuances in preferences. By adopting a more nuanced and inclusive framework that acknowledges and incorporates variations in response quality, we could extract richer insights and make more informed decisions regarding client preferences. Addressing these limitations could bolster the robustness and validity of our research, ultimately enhancing its relevance and impact in real-world applications.
Appendix A More Implementation Details
@@
A.1 Hyperparameter Settings
In our work, we fine-tune all models using LoRA, which is consistently set to rank 8, $\alpha=16$, and the dropout rate 0.0. For the generation, we apply with these parameters:
  • If it is required to generate multiple completions, then we set the temperature to 1.0.
@@ A.1 Hyperparameter Settings
SFT Selector Training RLFT
Participation Rate - 5/53 -
Local Iterations 30 30 30
Batch Size 32 16 32
Rounds 1000 500 500
Optimizer AdamW AdamW RMSprop
Hyperparameters (0.9, 0.95) (0.9, 0.95)
Learning rate $41e-$ 4 $51e-$ 5 $61e-$ 6
@@
Participation Rate 10/300 -
Local Iterations 10 10
Batch Size 16 16
Rounds 200 200
@@
Special Setting for FedBiscuit
@@
A.2 Instruction Tuning Prompt
In this section, we highlight the prompts used to fine-tune the summarization tasks and the QA task:
Summarization.
For pertaining (SFT) and the later reinforcement-learning fine-tuning (RLFT), it follows the prompt below
Below is a forum post. Write a precise and concise summary that includes the most important points of the post.
### SUBREDDIT: r/{subreddit}
### TITLE: {title}
### POST: {post}
### TL;DR:
For comparison:
Below is a forum post followed by two summaries. Pick a more precise and concise one that summarizes the most important points in the given forum post, without including unimportant or irrelevant details. State your choice with a single capital letter, i.e., Äïf SUMMARY A is better, B̈ïf SUMMARY B is better.
@@ A.2 Instruction Tuning Prompt
QA.
As the QA utilizes a pretrained model named Alpaca-7B, we follow its pretrained format
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
@@
For comparison between the two responses:
Below is a query followed by two responses. Pick a helpful response that is precise, concise, and casual. State your choice with a single capital letter, i.e., Äïf RESPONSE A is better, B̈ïf RESPONSE B is better.
@@
Appendix B More Numerical Results and Analysis
B.1 Numerical Results on QA
In this section, the test dataset comes from AlpacaFarm Dubois et al. (2024); Li et al. (2023b). The results are as follows:
2-completion 4-completion
Rating Rating
Alpaca-7B 3.752 -
FedBis 4.140 4.113
FedBiscuit 4.094 3.830
Table 4: Performce under QA.
As presented in Section 4, the LLM owner will generate a set of responses to a given instruction before building a preference dataset. Therefore, the column "2-completion" means the owner prepares 2 completions for each instruction, while "4-completion" means 4 completions for each instruction and forms 6 pairs. The row "Alpaca-7B" acts as a baseline to help us understand the performance of the proposed FedBiscuit and FedBis. All the rating comes from Auto-J Li et al. (2023a), which would be different from the ratings reported by Li et al. (2023b) because it evaluates with GPT-4 Achiam et al. (2023).
The table above may lead to conclusions different from those drawn from the summarization task. First, FedBis achieves better performance than FedBiscuit. This is within our expectations. First, these selectors are trained for a total of 200 rounds. As presented in Figure 3, FedBiscuit surpasses FedBis after 300 communication rounds. This is because the selectors of FedBiscuit are trained for 100 rounds only, while the selector of FedBis has been fully trained for 200 rounds. When the inflection point appears in FedBis, we can hypothesize that the dominance of FedBiscuit still exists.
Another comparison arises between different numbers of generations to a given prompt. From the table, we notice that "2-completion" can achieve better performance than "4-completion," meaning that the performance may not be relevant to the size of the RLFT training set. Instead, it may rely on the quality of the training data. As we can see, Alpaca-7B hardly generates high-quality data, leading to limited improvement with training with these generated data. In other words, if we generate more low-quality completions, the improvement of the model would be more limited. At the same time, we can hypothesize that if the generated data are of high quality, "4-completion" may outperform "2-completion" in terms of the final rating.
B.2 Examples on the Final RLHF Results
In this section, we prepare two examples to illustrate the performance of our proposed FedBis and FedBiscuit and compare them with other baselines. Refer to Table 5 and Table 6 for the demonstration of the examples.
  Subreddit: r/relationships Title: Me [25F] with my SO [27 M] of 3.5 years, I went through his texts and can’t stop thinking about something I saw. Post: I admit that going through his text messages was wrong and I do feel incredibly guilty about it now but the bigger issue for me is that I found something that I can’t stop thinking about. He was talking to his good female friend and I saw a text that said "am I being an idiot for leaving [my name]"? . I stopped reading after that because I started having a panic attack. The idea of breaking up was not even on my radar and I thought he was more than happy. Again, I know I shouldn’t have looked but whats done is done and now I can’t stop worrying about when it’s going to happen. And now i’m being clingy just because of my worry which i’m sure doesn’t help. I love him and don’t want to lose him.
Human Summary: I saw a text from my SO to his friend about leaving me, can’t stop freaking out. Need serious advice.
SFT: I went through my SO’s texts and found something I wasn’t supposed to. Now I’m being clingy and I don’t want to lose him.
Centralized: went through my boyfriends texts and found something that i’m now worrying about when/if he’s going to break up with me. I love him and don’t want to lose him. Help.
FedBis: went through SO’s texts and found a text that said "am I being an idot for leaving [my name]"? and now I’m worried he’s going to break up with me and can’t stop thinking about it.
FedBiscuit: went through SO’s texts, found something that i’m now worrying about when the inevitable breakup is going to happen. Now being clingy and can’t stop thinking about it. Help!
Table 5: An example for summarization task. Caution: The example data that may be offensive or harmful.
  List the layers of the TCP/IP model and for each layer identify one protocol.
Alpaca-7B: The TCP/IP model consists of four layers: the application layer, transport layer, internet layer and network access layer. Application Layer: HTTP
Transport Layer: TCP
Internet Layer: IP
Network Access Layer: Ethernet
FedBis: The TCP/IP model has four layers: the application layer, transport layer, internet layer and network access layer. At the application layer, the protocol is TCP. At the transport layer, the protocol is UDP. At the internet layer, the protocol is IP. At the network access layer, the protocol is Ethernet. 1. Application layer: TCP
2. Transport layer: UDP
3. Internet layer: IP
4. Network access layer: Ethernet
FedBiscuit: The TCP/IP model has four layers: the application layer, transport layer, internet layer and network access layer. At the application layer, the HTTP protocol is used. The transport layer uses the TCP protocol. At the internet layer, the IP protocol is used. Finally, the network access layer uses the Ethernet protocol.
Table 6: An example for QA task. Caution: The example data that may be offensive or harmful.

### v2

@@ A.2 Reinforcement-learning Fine-tuning
Appendix B Implementation Details and Hyperparameters
@@
B.1 Hyperparameter Settings
In our work, we fine-tune all models using LoRA, which is consistently set to rank 8, $\alpha=16$, and the dropout rate 0.05. For the generation, we apply these parameters:
  • If it is required to generate multiple completions, then we set the temperature to 0.7. We set the maximum new tokens for 80 under the summarization task and 300 for QA tasks.
@@ B.1 Hyperparameter Settings
Selector Training RLFT
Participation Rate 5/53 -
Local Iterations 30 -
Batch Size 32 32
Rounds 500 5 epochs
Optimizer AdamW RMSprop
Hyperparameters (0.9, 0.95) -
Learning rate $51e-$ 5 $61e-$ 6
@@
Participation Rate 10/200 -
Local Iterations 30 -
Batch Size 32 32
Rounds 500 5 epochs
@@
B.2 Dataset Details
In Section 6, we discuss how to partition the dataset for two tasks, namely, summarization and QA. Table 5 comprehensively presents the dataset details of both tasks, while Figure 3 visualizes the data distribution of the selected clients.
  Task Preference Dataset # preference samples # clients Partition Rules Max. Min. Std. Unlabelled Dataset # unlabelled prompts
Summarization TL;DR comparison 92858 53 Worker ID 12985 1 2284.33 Open-AI TL;DR 42782
Question Answering SHP 260814 200 Dirichlet(0.3) on categories 4393 260 832.39 SHP Test 4293
SafeRLHF 44578
Table 5: Dataset details for federated human preference benchmark
Special Setting for FedBiscuit.
@@ B.2 Dataset Details
B.3 More Experiments
  Methods Win Rate on Qwen-2 (0.5B) (# Wins / # Ties) Win Rate on Qwen-2 (1.5B) (# Wins / # Ties)
FedBis 86.69% (5681 / 26) 82.45% (5401 / 34)
FedBiscuit ($U=3$) 78.45% (5141 / 29) 80.18% (5254 / 29)
FedBiscuit ($U=5$) 73.97% (4847 / 22) 76.22% (4995 / 43)
Table 6: Performance using Gemma-2B to summarize a Reddit post (i.e., summarization task). Bold means the best result under the selector.
Comparison between two different model sizes.
Table 6 compares the performance between two Qwen-2 models as the selectors with different sizes (i.e., 0.5B and 1.5B) when using them to fine-tune a Gemma-2B model for the summarization task. These results show that while larger binary selectors (e.g., Qwen-2-1.5B) sometimes provide slight performance improvements, smaller selectors like Qwen-2-0.5B remain competitive, particularly when applying FedBis. These findings suggest that both model type and size influence the selector’s effectiveness, and they need to be carefully balanced based on task requirements and resource constraints.
B.4 Instruction Tuning Prompt
The proposed work follows a previous study Zhang et al. (2024) to fine-tune the model following a given prompt template. This is also known as instruction tuning. The prompt template is different between tasks and between the selector and base model. As a result, we provide the prompts in Figure 4 for detailed study.
Prompts template of LLM generation for summarization task. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
Summarize the following Reddit post in a paragraph of 50 words or less.
### Input:
SUBREDDIT: r/{subreddit}
TITLE: {title}
POST: {post}
### Response:
Prompt template when training a binary selector for summarization task. Below is a forum post followed by two summaries. Pick a more precise and concise one that summarizes the most important points in the given forum post, without including unimportant or irrelevant details. State your choice with a single capital letter, i.e., “A” if SUMMARY A is better, “B” if SUMMARY B is better.
@@ B.4 Instruction Tuning Prompt
Prompts template of LLM generation for QA task with an additional input. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
@@
Prompts template of LLM generation for QA task without an additional input. Below is an instruction that describes a task. Write a response that appropriately completes the request.
### Instruction:
{instruction}
### Response:
Prompt template when training a binary selector for QA task with SHP dataset. Below is a query followed by two responses. Pick a helpful response that is precise, concise, and casual. State your choice with a single capital letter, i.e., “A” if RESPONSE A is better, “B” if RESPONSE B is better.
@@
Figure 4: Prompt Templates.
Appendix C Sample Dialogs
We provide sample dialogs for our methods, FedBiscuit and FedBis, compared against the baselines FedDPO, FedAvg, and the raw model, highlighting their performance in summarization and question-answering tasks. Our methods consistently deliver superior results, as demonstrated by the provided samples.
FedBiscuit generates precise and well-organized summaries, making it particularly effective in distilling complex information into clear and actionable insights. FedBis, while maintaining a similar level of clarity, offers more nuanced and in-depth responses, making it versatile for a wider range of queries. In contrast, the baseline models—FedDPO, FedAvg, and the raw model—tend to produce less coherent responses. FedDPO sometimes lacks structure, FedAvg oversimplifies the content, and the raw model can miss critical details, leading to less accurate results.
These samples provide clear evidence that FedBiscuit and FedBis offer better summarization and question-answering performance by consistently striking the right balance between conciseness, accuracy, and detail, proving to be more reliable and contextually appropriate than the baselines.
User Input Give me a brief scenario of a persona that would search this information and find this content helpful: {NSW Residential Building Contract for Small Works over \$20,000 Designed for residential building work which is not complex over \$20,000. Suitable for smaller alteration and renovation projects. Small Works Contracts (Form 10A) Pricing structure Fixed price Value range Under \$7,500 Use for Minor residential alterations and additions work Includes Succinct, plain-language conditions suitable for very small, low-risk projects in place of a quotation}
FedBiscuit Imagine you are a homeowner who wants to make some changes to your home, but you don’t want to spend a lot of money on a major renovation. You come across a small works contract that is designed for residential building work that is not complex and has a value range of under \$7,500. This contract is suitable for smaller alteration and renovation projects. It is a fixed price contract, meaning the price is set and there are no hidden costs or additional fees. The contract is written in plain language, making it easy to understand and use. It includes succinct conditions that are suitable for very small, low-risk projects. This means that if something goes wrong, you can easily resolve the issue and the contract will protect you. Overall, this small works contract for residential building work is a great option for homeowners who want to make some changes to their home without spending a lot of money. It is a simple, straightforward solution that provides peace of mind and protection for your project.
FedBis Imagine you are a homeowner who wants to make some changes to your home, but you don’t want to spend a lot of money on a major renovation. You come across a small works contract that is designed for residential building work that is not complex and has a value range of under \$7,500. This contract is suitable for smaller alteration and renovation projects. It is a fixed price contract, meaning the price is set and there are no hidden costs or additional fees. The contract is written in plain language, making it easy to understand and use. It includes succinct conditions, which are suitable for very small, low-risk projects. The contract is intended for residential work, meaning it is not suitable for larger, more complex projects. Overall, this small works contract for residential building work is a cost-effective and straightforward option for making small changes to your home without breaking the bank.
FedDPO A persona that would find this information helpful is a homeowner who is planning to make some minor changes to their home, such as adding a new room or making some renovations. They are looking for a simple and straightforward contract that outlines the terms and conditions of the work, with a fixed price and a value range of under \$7,500. They want to ensure that the contract is suitable for their small, low-risk project and that it is concise and easy to understand.
FedAvg I’m not sure if this is helpful, but I’m a lawyer and I’ve been looking for a similar contract for a while now. I’ve found a few, but they’re not quite what I’m looking for.
Raw Model A persona that would search this information and find it helpful is a small residential building contractor who is looking to complete a small works project on their home. They are looking for a fixed price contract that is suitable for their project, which is not complex and will not exceed \$20,000. They want a contract that is simple and easy to understand, with concise conditions that outline the scope of work and the expected timeline. They are looking for a contract that is suitable for small, low-risk projects, and they want to avoid the need for a quotation.
Figure 5: Gemma QA Sample.
User Input Hi, I have some falafel, but no tahini to put on them. Can you give me a recipe for making tahini?
FedBiscuit Sure, here is a recipe for making tahini at home: Ingredients: - 1 cup of sesame seeds - 2 tablespoons of water - 2 tablespoons of honey - 2 tablespoons of lemon juice Instructions: 1. Place the sesame seeds in a food processor and process until they become a paste. 2. Add the water, honey, and lemon juice and process until the mixture is smooth. 3. Taste and adjust the seasoning if necessary. 4. Transfer the tahini to an airtight container and store it in the refrigerator for up to 1 month.
FedBis Sure, here is a recipe for making tahini at home: Ingredients: - 1 cup of sesame seeds - 2 tablespoons of water - 2 tablespoons of honey - 2 tablespoons of lemon juice - 1 teaspoon of garlic - 1/2 teaspoon of salt Instructions: 1. Place the sesame seeds in a food processor and process until they become a paste. 2. Add the water, honey, lemon juice, garlic, and salt and process until the mixture is smooth. 3. Taste and adjust the seasoning if necessary. 4. Transfer the tahini to an airtight container and store it in the refrigerator for up to 1 month.
FedDPO You can make tahini at home by soaking a few tablespoons of sesame seeds in a small amount of water for 30 minutes, then grinding them in a food processor. Add a tablespoon of the resulting paste to your falafel to give them a delicious, nutty flavor.
FedAvg I’ve been making tahini with a blender. I put a cup of sesame seeds in a blender, add a cup of water, and blend until smooth. I then strain the mixture through a cheesecloth-lined strainer. I’ve been using the tahini for a few weeks now and it’s been great.
Raw Model You can make tahini at home by soaking a few tablespoons of sesame seeds in a small amount of water for 30 minutes, then grinding them in a food processor. Add a tablespoon of the resulting paste to your falafel to give them a nutty flavor.
Figure 6: Llama QA Sample.
Reddit Post: What is your most memorable act of kindness and why? Random or otherwise. I was walking to school one morning when I was about 15 when I heard a voice call out "Hello? Excuse me? Young man?". I turned and saw a frail old man who must have been in his 90s standing in the doorway of his house. I walked over and he asked if I could help him with something. I said I would help, so I followed him to a step-ladder in his kitchen. He said that he needed to change the bulb but he couldn’t get up the ladder. After I changed the bulb, he said that he’d give me something for helping out. He reached for a small tray of change on his kitchen table and picked out the only 50p coin in his tray of 1p and 2p coins. I told him he didn’t have to give me anything but he said that it was only right to reward helping out an old man.
FedBiscuit A Reddit user recounts a heartwarming act of kindness where they helped an elderly man change a lightbulb in his kitchen. The stranger gave them a 50p coin as a reward for their help. The act of kindness was memorable because it brought a smile to the elderly man’s face and showed that even in their old age, people can still do good deeds for others.
FedBis A Reddit user recounts a heartwarming act of kindness where they helped an elderly man change a lightbulb in his kitchen. The stranger gave them a 50p coin as a reward for their help. The act of kindness was memorable because it brought a smile to the elderly man’s face and showed that even in their old age, people can still do good deeds for others.
FedDPO A Reddit user recounts a heartwarming story of a 90-year-old man who needed help changing a lightbulb in his kitchen. The user, who was 15 at the time, heard the man’s plea for help and went to his aid. The man gave the user
FedAvg Old man in his 90s needed help changing a bulb. I helped him and he gave me a 50p coin.
Raw Model A Reddit user recounts a heartwarming story of a stranger who helped an elderly man change a lightbulb in his kitchen. The stranger, who was walking to school at 15, heard a voice call out for help and followed the old man to his house. He changed the lightbulb and the old man gave him a 50p coin as a reward. The stranger didn’t have
Figure 7: Gemma Summarization Sample.
Reddit Post: TIFU by locking myself out of my room. It all started with a cultural tour. Being active duty military and stationed overseas in Japan we were voluntold to partake in a tour of the local area. Fun and dandy right? I was ecstatic about the free tour of the local area. Except when I woke up five minutes before the tour starts. Luckily the meeting area was a Starbucks just outside of my room. Que the rushing, the frantic panic to dress up, and get down in time. I left my ID which is also conveniently my room key. Two more mistakes. (Straying away from the story, the military is all about being on time and integrity. Hence "excellence in all we do".) Realizing I fucked up bad, I told the host that I had left my REQUIRED ID in my room, which is my room key locking myself out. I couldn’t go on the tour and now I’m stuck in the Starbucks waiting until one of my supervisors picks me up and lets me into my room. I’m expecting more than just a little scolding, paperwork even.
FedBiscuit The poster was locked out of their room due to leaving their ID, which was also their room key, in their rush to make it on time for a cultural tour. They are now stuck in a Starbucks waiting for someone to pick them up and let them into their room. They expect to receive a scolding and possibly even paperwork for the mistake.
FedBis While on a cultural tour in Japan, the poster locked themselves out of their room, leading to them being stuck in a Starbucks and expecting a scolding from their supervisors.
FedDPO While on a cultural tour in Japan, the poster was locked out of their room due to leaving their ID and room key in the room. They are now stuck in a Starbucks waiting for a supervisor to pick them up and let them into their room. They expect to receive a scolding and possibly paperwork for the mistake.
FedAvg I locked myself out of my room and now I’m stuck in Starbucks waiting for a supervisor to pick me up and let me into my room.
Raw Model While on a cultural tour in Japan, the poster was locked out of their room due to leaving their ID and room key in the room. They were stuck in a Starbucks until one of their supervisors picked them up and let them into their room. They expect to receive a scolding and possibly paperwork for the mistake.
Figure 8: Llama Summarization Sample.